Hi all, I was hoing to get some help with the following:
I need to make t the subject of the following expression:
h^3 = ln(g /(t^2+1))
Help with steps and explanations would be great.
Thanks, -Nick.

Answer by stanbon(75887)

make t the subject of the following expression:
h^3 = ln(g /(t^2+1))
e^(h^3) = g/(t^2+1)
t^2+1 = g/e^(h^3)
t = sqrt{[g/e^(h^3)]-1}
============================
Cheers,
Stan H.